Rosaline consumes only turkey sandwiches and juice and she spends all her income on turkey sandwiches and juice. The marginal utility of her last turkey sandwich consumed is 200 utils and the marginal utility of her last juice consumed is 100 utils. The price of a turkey sandwich is $4 and the price of juice is $3.
1. Is Rosaline maximizing her level of utility? Justify.
1
Expert's answer
2020-05-13T11:07:47-0400

The utility refers to the usefulness of a commodity or the satisfaction that is received by an individual as a result of consuming a product.

An individual will maximize the utility under the following condition:

"\dfrac{\text{Mu}_\text{t}}{\text{Mu}_\text{j}}=\dfrac{\text{P}_\text{t}}{\text{P}_\text{j}}"


"\text{Mu}_\text{t}" is the marginal utility of Turkey sandwich.

"\text{Mu}_\text{j}" is the marginal utility of juice.

"\text{P}_\text{t}" price of Turkey sandwich.

"\text{P}_\text{j}" price of juice.

Substituting the given information to the above condition we get:

"\dfrac{200}{100}=\dfrac{4}{3}"

"2>\dfrac{4}{3}"

The utility maximizing condition is not satisfied, therefore, the utility is not maximized.